solve the equation. If the exact roots cannot be found, state the consecutiveintegers between which roots are located.
-2x^2-4x=0
I don't not understand any part of this. I don't even know where to begin. I would greatly appreciate a step by step solution so I can understand how to do the rest of the problems.

Answer by edjones(8007)

-2x^2-4x=0
-2x(x+2)=0
x=0, x=-2
.
Ed